OpenStudy (anonymous):

A medical scientist has a 15-gram sample of I-131 and would like to its half-life, in days. He also knows that k = 0.0856. N = N0e -kt N0 = 15 grams N = 7.5 grams k = 0.0856 t = ? Find the half-life, in days, of I-131 using the information at the left. Round your answer to the nearest tenth.

OpenStudy (photon336):

Please re-arrange the formula for t \[N = N_{0}e^{-kt}\] \[\frac{ N }{ N_{0} } = e^{-kt}\] \[\ln(\frac{ N }{ N_{0} }) = \ln(e^{-kt})\] \[Ln(\frac{ N }{ N_{0} }) = -kt \] \[-k*\ln (\frac{ N }{ N_{0} }) = t \]

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

And don't I need to divide 7.5 abd 0.0856

OpenStudy (anonymous):

And*

OpenStudy (photon336):

\[-8.56*10^{-2}*\ln (\frac{ 7.5 }{ 15 }) = t\]